[figure: 9.7 cm]

SUBMILLIMETER SPECTROSCOPY OF ANTIFERROMAGNETIC FERROELECTRIC YMn_{2}O_{5}

A.A. Mukhin ^{a}, V.D. Travkin ^{a}, V.Yu. Ivanov ^{a}, S.P. Lebedev, A.S. Prokhorov ^{a} and K. Kohn ^{b} aGeneral Physics Institute of the Russian Acad. Sci., 38 Vavilov St., 117942 Moscow, Russia b Waseda University, Shinjuku-ku, Tokio 169, Japan

Dynamics properties of the antiferromagnetic ferroelectric Y Mn_{2}O_{5} were studied in the millimeter-submillimeter frequency range (2-33 cm^{−1}) in temperatures 5-300 K. Four absorption lines were observed in transmission spectra. Three of them appeared as a wide absorption band at 15-25 cm^{−1} below T_{N}=40 K, then after antiferromagnetic helical ordering of Mn ions, it splits into three lines. The fourth, most intensive line appears at the lower frequencies (6-8 cm^{−1}) below 20 K after a second phase transition. Temperature dependence of resonance frequencies, linewidths and mode contributions to magnetic and dielectric susceptibilities were obtained. The fourth most intensive mode was found to have dielectric origin. Three high frequency modes are probably due to an antiferromagnetic (exchange) resonance, excited both by ac magnetic and electric fields. The work was supported in part by the RFBR 98-02-17740.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Alexander A. Mukhin

Address for correspondence : General Physics Institute of the Russian Acad. Sci., Vavilov St., 38. 117942 Moscow, Russia

Email address : mukhin@ran.gpi.ru